OpenStudy (anonymous):

simplifying factorials: (2k+1)!/(2k+3)!

OpenStudy (anonymous):

write it out, it will be obvious really, try it

OpenStudy (anonymous):

You have n! =n * (n-1)! so 3! = 3*2*1. 4! / 2! = 4*3*2/2 = 4*3, etc.

OpenStudy (perl):